OpenStudy (anonymous):

Write another possible formula for y = −2 cos(3x + π) + 1 using the sine function and a positive coefficient on the sine.

myininaya (myininaya):

remember \[\cos(x)=\sin(\frac{\pi}{2}-x)\] \[y=-2\sin(\frac{\pi}{2}-(3x+\pi))+1\]

myininaya (myininaya):

you can simplify that inside

OpenStudy (anonymous):

wait, so you have t plug it in? so how would you handle the pi in the original equation? ugh I'm really bad at all these identities.

myininaya (myininaya):

you have \[y=-2\sin(\frac{\pi}{2}-3x-\pi)+1=-2\sin(\frac{\pi}{2}-\frac{2\pi}{2}-3x)+1\] \[y=-2\sin(\frac{-\pi}{2}-3x)+1\] and it also say to write with positive coeifficient so also recall that sin(-x)=-sinx so we have \[y=-2\sin(-[\frac{\pi}{2}+3x])+1=-2*-\sin(\frac{\pi}{2}+3x)+1=2\sin(\frac{\pi}{2}+3x)+1\]